# Nonlinear optimals and their role in sustaining turbulence in channel flow

### v1

@@ 1 Introduction
A natural generalisation, which also helps address these challenges, is to consider the full nonlinear optimisation problem. Thanks to the presence of the nonlinear term, the primary and secondary linear processes can be coupled together, thus eliminating the need for formulating a two-stage process. Of course, it is unclear a priori if the nonlinear optimal actually involves such a two-stage process, or if other dynamics are more effective. More generally, the central question to be addressed in this work is which aspects of the nonlinear optimal resemble the dynamics of real turbulence.
@@
By comparison, the nonlinear optimisation problem has received much less attention. For one, nonlinear optimisation relying on solving the full three-dimensional Navier–Stokes equations only became computationally feasible during the 2010s (Pringle & Kerswell, 2010; Cherubini et al., 2010, 2011; Monokrousos et al., 2011; Pringle et al., 2012). However, even today, the computational effort remains high, leading to relatively few works applying nonlinear optimisation to the turbulent setting (Cherubini & De Palma, 2013; Farano et al., 2015, 2017). Notably, however, these works employ a different formulation than used here, rendering these works unsuitable to answer the questions we are presently concerned with. This is discussed in greater detail in Section 2.
@@ 2 Problem formulation
$\mathcal{N}{}$ :-$\coloneq\cfrac{\partial\mathbf{U}}{\partial t}+\mathbf{U}\cdot\nabla\mathbf{U}+\nabla P-\frac{1}{\text{Re}_{\tau}}\nabla^{2}{\mathbf{U}}=0,:$ (1)
@@
where $\mathbf{U}$ and $P$ are the dimensionless flow velocity and pressure, respectively. The coordinate system is oriented such that $x$ points in streamwise direction, $y$ in wall-normal direction and $z$ in spanwise direction. Decomposing velocity and pressure into a base and a disturbance part,
$\mathbf{U}=\bar{\mathbf{U}}+\mathbf{\tilde{u}},\quad P=\bar{P}+\tilde{p},$ (3)
@@
$\overline{\mathcal{N}}{}$ :-$\coloneq\overline{\mathbf{\tilde{u}}\cdot\nabla\mathbf{\tilde{u}}}-\mathbf{\hat{x}}-\frac{1}{\text{Re}_{\tau}}\cfrac{d^{2}\bar{\mathbf{U}}}{dy^{2}}=0,:$ (4)
$\mathcal{M}{}$ :-$\coloneq\nabla\cdot\bar{\mathbf{U}}{}=0,:$ (5)
where we have used that $\bar{\mathbf{U}}=\bar{\mathbf{U}}(y)$ and $\bar{U}_{y}=\bar{U}_{z}=\mathbf{\bar{\mathbf{\tilde{u}}}}=0$. We can derive an equation for the disturbance fields by subtracting Eqs. 4 and 5, from Eqs. 1 and 2, yielding
$\tilde{\mathcal{N}}{}_{\text{full}}$ :-$\coloneq\cfrac{\partial\mathbf{\tilde{u}}}{\partial t}+\bar{\mathbf{U}}\cdot\nabla\mathbf{\tilde{u}}+\mathbf{\tilde{u}}\cdot\nabla\bar{\mathbf{U}}+\mathbf{\tilde{u}}\cdot\nabla\mathbf{\tilde{u}}-\overline{\mathbf{\tilde{u}}\cdot\nabla\mathbf{\tilde{u}}}+\nabla\tilde{p}-\frac{1}{\text{Re}_{\tau}}\nabla^{2}{\mathbf{\tilde{u}}}=0,:$ (6)
@@
By further fixing the initial energy $e_{0}$, one can ask which initial condition $\mathbf{\tilde{u}}(\mathbf{x},0)$ experiences the highest energy gain $G$ by the end of the time horizon $T$. This question naturally leads to an optimisation problem which has been studied extensively in previous works (Pringle & Kerswell, 2010; Cherubini et al., 2010, 2011; Monokrousos et al., 2011; Cherubini & De Palma, 2013; Farano et al., 2017), and which can be solved using adjoint looping. Details of the solution algorithm are given in Section 3.
A subtle but important point must be made about the role of the Reynolds stress tensor $\overline{\mathbf{\tilde{u}}\cdot\nabla\mathbf{\tilde{u}}}$, which appears in 6 because $\bar{\mathbf{U}}$ is the temporal mean and not a solution of the Navier–Stokes equations 1. This term ensures that the disturbances $\mathbf{\tilde{u}}$ are of the appropriate magnitude to produce a Reynolds stress tensor capable of maintaining $\bar{\mathbf{U}}$. However, for questions the present investigation is concerned with, this particular feedback mechanism must be suppressed, as is discussed below. Therefore, instead of Eq. 6, we consider the modified momentum equation without the Reynolds stress tensor $\overline{\mathbf{\tilde{u}}\cdot\nabla\mathbf{\tilde{u}}}$,
$\tilde{\mathcal{N}}{}\coloneq\cfrac{\partial\mathbf{\tilde{u}}}{\partial t}+\bar{\mathbf{U}}\cdot\nabla\mathbf{\tilde{u}}+\mathbf{\tilde{u}}\cdot\nabla\bar{\mathbf{U}}+\mathbf{\tilde{u}}\cdot\nabla\mathbf{\tilde{u}}+\nabla\tilde{p}-\frac{1}{\text{Re}_{\tau}}\nabla^{2}{\mathbf{\tilde{u}}}=0.$ (8)
In this study, only Eqs. 8 and 7 are actually solved numerically. The base velocity is imposed to be the time-averaged mean profile obtained from a previous DNS (Markeviciute & Kerswell, 2024).
Considering Eq. 8 instead of Eq. 6 crucially differentiates the present investigation from previous works such as Farano et al. (2017), who do include the Reynolds stress tensor in their evolution equation. Note that Eq. 8 effectively implies that the base flow is maintained by an external forcing, thus “releasing” the turbulent velocity disturbances from their responsibility of maintaining the base flow, and allowing them to develop freely. On the other hand, if the Reynolds stress tensor is included, the disturbance energy is forced to be at a level consistent with fully-developed turbulence within the first few time steps of the simulation. Whether this is desired strongly depends on the question being investigated. In the present study, there are two reasons why this would be inappropriate: First, we would like to consider the initial disturbance energy as a parameter and also study much smaller disturbances than would be found in fully-developed turbulence, much in the spirit of a minimal seed calculation. With the Reynolds stress tensor fixing the order of magnitude of the final disturbance energy, however, the gain would no longer be a good measure to compare vastly different initial disturbance energies. Second, we are interested in the energy transfer from the mean to the fluctuations, the isolating of which requires removing the Reynolds stress tensor from the equations.
@@
where $\lambda,\mathbf{\tilde{v}}$ and $\tilde{q}$ are Lagrange multipliers. Setting the first variation of $\mathcal{L}$ to zero leads, after some algebra, to the conditions
@@
$\cfrac{\partial\mathbf{\tilde{v}}}{\partial t}+(\bar{\mathbf{U}}+\mathbf{\tilde{u}})\cdot\nabla\mathbf{\tilde{v}}-\mathbf{\tilde{v}}\cdot{(\nabla(\bar{\mathbf{U}}+\mathbf{\tilde{u}}))}^{\intercal}+\nabla\tilde{q}+\frac{1}{\text{Re}_{\tau}}\nabla^{2}{\mathbf{\tilde{v}}}$ $=0,$ (11)
@@
Note that Eqs. 12 and 11 form the adjoint Navier–Stokes equations, and the Lagrange multipliers $\mathbf{\tilde{v}}$ and $\tilde{q}$ can be interpreted as adjoint velocity and pressure (divided by the density), respectively. This set of equations naturally gives rise to a looping algorithm to obtain the optimal initial condition $\mathbf{\tilde{u}_{\text{opt}}}(\mathbf{x},0)$ that generates the highest energy gain $G$ at the end of the time horizon $T$: First, make an initial guess for $\mathbf{\tilde{u}_{\text{opt}}}(\mathbf{x},0)$, which we call $\mathbf{\tilde{u}}(\mathbf{x},0)$, and evolve it forward in time using the Navier–Stokes equations 7 and 8 until time $t=T$. Second, obtain the initial condition for the adjoint system $\mathbf{\tilde{v}}(\mathbf{x},T)$ using Eq. 10 and integrate Eqs. 12 and 11 backward in time from $t=T$ to $t=0$. Third, now that the final value of the adjoint variable $\mathbf{\tilde{v}}(\mathbf{x},0)$ is known, calculate ${\delta\mathcal{L}}/{\delta\mathbf{\tilde{u}}(\mathbf{x},0)}$ using Eq. 13 and use this gradient to improve the current guess $\mathbf{\tilde{u}}(\mathbf{x},0)$. Repeat this loop until some convergence criterion is reached. More technical details of this algorithm are discussed in Section 3.
@@ 3 Computational method
A special-purpose solver has been implemented in the programming language Python to integrate 7 and 8 forward in time and to integrate the adjoint system 12 and 11 backward in time. This solver utilises the library JAX (Bradbury et al., 2018), which makes the code fully differentiable (although this is not used much in the present work) and provides GPU-acceleration, thus delivering a crucial speed-up and making the investigation feasible. The source code of the solver is available at https://github.com/dakling/jax-spectral-dns.
@@
The effectiveness and efficiency of the looping approach crucially depends on the employed gradient-based optimisation algorithm. In the present work, we follow Cherubini et al. (2011) and use a conjugate gradient method, although we determine the step size adaptively using a two-way backtracking line search (Truong & Nguyen, 2020). This line search is similar to a regular backtracking line search, but the step size from the last iteration is used as an initial guess for the new step size, thus reducing the number of necessary function evaluations (which are quite expensive as they involve the forward run of the DNS). If the step size is found to be sufficiently small, it is also periodically tested if the step size can be increased, which is not necessary in classical backtracking line search. Note that the formula for the gradient 13 contains the Lagrange multiplier $\lambda$, which ensures that the initial energy constraint is fulfilled. Its appropriate value depends on the step size and direction taken, so the gradient itself depends on the parameters of the conjugate descent and must be recomputed appropriately. For a given set of descent parameters, $\lambda$ is then obtained using a standard Newton method. To ensure that numerical issues with determining $\lambda$ can never lead to a violation of the initial energy constraint, the resulting initial condition is also appropriately rescaled after each optimisation iteration. The algorithm is documented in detail in Section A.3.
@@ 4 Results
For any $T$, as $e_{0}\rightarrow 0$, the linear optimisation problem is recovered. Consequently, for small $e_{0}$, a quasilinear optimal, which closely resembles the linear optimal, is obtained. As $e_{0}$ increases, there comes a point where a structurally completely different nonlinear optimal arises. For large $e_{0}$, the final state of the disturbances may be well in the turbulent regime, which, in the current formulation, usually leads the optimisation algorithm to fail, as the final state dependency on the initial condition is too sensitive and chaotic for the optimisation problem to remain well-conditioned. The precise determination of this boundary is difficult and not very important to the present study, but values up to $e_{0}/E_{0}=10^{-4}$ were found to converge without issues even for the longest time horizons considered here, where $E_{0}$ denotes the energy of the base flow. Naturally, we are most interested in nonlinear optimals whose initial energy is below this value.
@@
Figure 1: Optimal gain over initial energy normalised by the energy of the base flow $E_{0}$. Note the truncated vertical axis. Plots of the initial condition of the streamwise disturbance velocity for some of the optimals are also included. For more details on these and their time evolution please refer to Figs. 3 and 4.
It is apparent that at $e_{0}/E_{0}={2}\cdot 10^{-5}$, a first nonlinear optimal emerges. As the initial energy increases further, this optimal becomes more and more efficient, until the highest gain among all $e_{0}$ is found at $e_{0}/E_{0}={7.2}\cdot 10^{-5}$. As $e_{0}/E_{0}$ is increased beyond this value, saturation seems to set in, and the achieved gains slowly diminish, until around $e_{0}/E_{0}={2}\cdot 10^{-4}$, the optimisation problem starts becoming ill-conditioned. Interestingly, the difference between the highest nonlinear optimal gain ($34.82$) and the linear gain ($28.48$) is not very high at this $T$. However, the long-term evolution reveals a major difference: whereas the linear optimal tends to reach its peak gain at or close to $t=T$, the nonlinear optimals keep growing significantly, often almost by another order of magnitude (see Fig. 2).
Figure 2: Long-term evolution ($t_{\text{final}}u_{\tau}/h=60$) of the energy gain of the quasilinear optimal ($e_{0}/E_{0}=10^{-6}$) and the nonlinear optimal ($e_{0}/E_{0}={7.2}\cdot 10^{-5}$), both for $Tu_{\tau}/h=0.7$ (dashed vertical line).
@@
Moving to finite but small values for $e_{0}/E_{0}$, at $10^{-6}$ and up to $10^{-5}$, a quasilinear optimal is found. Its velocity time evolution (not shown) closely resembles the linear optimal (Fig. 3), though the effects of the nonlinear term are visible in the combined energy of the non-dominant wave modes, which grows from around ${3}\cdot 10^{-4}$ to ${2.6}\cdot 10^{-3}$ times the total energy.
Optimal gain is achieved by the nonlinear optimal with initial energy $e_{0}/E_{0}={7.2}\cdot 10^{-5}$, and its time evolution is shown in Figs. 4 and 5. It is apparent that the shape of the initial condition is completely different from the linear optimal. Most notably, no streaks are present, and the disturbance is concentrated, or localised, in specific parts of the domain. This localisation is most clear in spanwise direction, as there is no full localisation in the streamwise direction (see Fig. 5). Instead, in streamwise direction, there seems to be a length scale between the two distinct peaks of the disturbance. Indeed, when moving to a box twice as large in streamwise direction (i.e. $2\pi\times 2\times\pi$), four such peaks can be identified (see Section A.2). It is unclear how this length scale is determined, but we may speculate that it is the most efficient spacing to quickly generate continuous streaks. This may also be related to the size of the minimal channel unit, which is roughly half of the current domain in streamwise direction. Therefore, it is possible that the current structure could be thought of as two minimal channel optimals stitched together in streamwise direction. Localisation is a well-known phenomenon for nonlinear optimals (Kerswell, 2018), and can be rationalised by the optimisation algorithm trying to accommodate the initial energy constraint: instead of distributing the disturbance energy evenly, it may make more sense to have high-amplitude disturbances strategically placed in some small areas and then rely on nonlinear mechanisms to spread out these disturbances over time.
@@
The early part of the evolution is characterised by an absence of streaks, and instead we observe tilted flow structures oriented such that they are amplified by an Orr-type mechanism (this is most clearly observed in Fig. 7). Following this initial phase, these disturbances quickly organise into streaks (see Figs. 5 and 6, middle tile). Note that the streaks that have formed at this point do exhibit correct streak spacing and amplitude (see Fig. 4, middle tile, and the discussion in Section 4.2). Unlike the streaks of the linear optimal, the low-speed streaks do not remain straight as they evolve, but meander and wrap around the high-speed streaks, as is revealed by Fig. 6. Subsequently, these streaks grow due to the lift-up mechanism, and although no clear streak breakdown can be observed, it is clear from Fig. 8 that second-order disturbances grow on top of the streaks towards the end of the considered time window (from time (c) onwards). Thus, it can be concluded that the primary and, although not as pronounced, the secondary linear process are both visible in the nonlinear optimum, though the way that they are coupled together by the nonlinear term means that there is significant temporal overlap. This serves as a reminder that characterising them as neatly separated subprocesses constitutes a major simplification. Moreover, the streak growth itself is clearly not exclusively fuelled by transient growth (although this likely plays a significant role), as can be most clearly observed by the temporal evolution of the high-frequency modes ($k_{x}\geq 6$) between the times (a) and (b) shown in Fig. 8. After an initial growth up to the time marked by vertical line (a), these modes start decaying, indicating that they transfer energy to the streaks.

### v2

@@ 1 Introduction
A natural generalisation, which also helps address these challenges, is to consider the full nonlinear optimisation problem. Thanks to the presence of the nonlinear term, the primary and secondary linear processes can be coupled together, thus eliminating the need for formulating a two-stage process. Of course, it is unclear a priori if the nonlinear optimal actually involves such a two-stage process, or if completely different dynamics are more effective. More generally, the central question to be addressed in this work is which aspects of the nonlinear optimal resemble the dynamics of real turbulence.
@@
By comparison, the nonlinear optimisation problem has received much less attention. For one, nonlinear optimisation relying on solving the full three-dimensional Navier–Stokes equations only became computationally feasible during the 2010s (Pringle & Kerswell, 2010; Cherubini et al., 2010, 2011; Monokrousos et al., 2011; Pringle et al., 2012; Farano et al., 2016). However, even today, the computational effort remains high, leading to relatively few works applying nonlinear optimisation to the turbulent setting (Cherubini & De Palma, 2013; Farano et al., 2015, 2017). Notably, however, these works employ a different formulation than used here, rendering these works unsuitable to answer the questions we are presently concerned with. This is discussed in greater detail in Section 2.
@@ 2 Problem formulation
$\mathcal{N}{}$ :-$\coloneq\cfrac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla\mathbf{u}+\nabla p+\mathbf{f}-\frac{1}{\text{Re}_{\tau}}\nabla^{2}{\mathbf{u}}=0,:$ (1)
@@
where $\mathbf{u}$ and $p$ are the dimensionless flow velocity and pressure, respectively, and $\mathbf{f}$ is an external volume forcing term, the reasons for including which will become apparent later. The coordinate system is oriented such that $x$ points in streamwise direction, $y$ in wall-normal direction and $z$ in spanwise direction. We decompose velocity and pressure into a base and a disturbance part,
$\mathbf{u}=\mathbf{U}+\mathbf{\tilde{u}},\quad p={P}+\tilde{p},$ (3)
where capital letters denote the base and tilde denotes the disturbance fields. In this study, the base is obtained by temporal averaging an a priori DNS run. Since it is calculated in advance, $\mathbf{U}$ can be thought of as a fixed parameter entering the problem. In the following, we set $-\nabla{P}=\mathbf{\hat{x}}$, with $\mathbf{\hat{x}}$ denoting the unit vector in $x$-direction. Note that the streamwise-spanwise (but not temporal) average of $\mathbf{\tilde{u}}$, which we denote $\langle{\mathbf{\tilde{u}}}\rangle_{xz}$, is allowed to be non-zero. Here, angular brackets denote averaging across the directions in the index, so, for example,
@@
This ensures consistency because $\mathbf{\tilde{u}}$ is able to change the streamwise-spanwise averaged velocity $\langle{\mathbf{u}}\rangle_{xz}$ as the disturbance grows on the short time scales we are presently interested in, even though $\mathbf{U}$ remains fixed. However, in the limit of very long time horizons, the temporal average $\langle{\mathbf{\tilde{u}}}\rangle_{t}$ vanishes. Inserting Eq. 3 into Eqs. 2 and 1 yields
[…]
$+$ $\cfrac{\partial\mathbf{\tilde{u}}}{\partial t}+\mathbf{U}\cdot\nabla\mathbf{\tilde{u}}+\mathbf{\tilde{u}}\cdot\nabla\mathbf{U}+\mathbf{\tilde{u}}\cdot\nabla\mathbf{\tilde{u}}+\nabla\tilde{p}-\frac{1}{\text{Re}_{\tau}}\nabla^{2}{\mathbf{\tilde{u}}}=0,$ (5)
$\mathcal{M}{}\coloneq$ :- $\nabla\cdot\mathbf{U}{}+\nabla\cdot\mathbf{\tilde{u}}{}=0.$ (6)
We now set
$\mathbf{f}=\mathbf{\hat{x}}+\frac{1}{\text{Re}_{\tau}}\nabla^{2}{\mathbf{U}}.$ (7)
Physically, this forcing term represents the background turbulence fluctuations maintaining $\mathbf{U}$. It is important to note that these background fluctuations are separate from the disturbances $\mathbf{\tilde{u}}$ in our present model. This point is further discussed below. Together with ${\partial\mathbf{U}}/{\partial t}=0$ due to statistical stationarity and $\mathbf{U}\cdot\nabla\mathbf{U}=\nabla\cdot\mathbf{U}=0$ in channel flow, this yields an equation system for the disturbances which reads
$\tilde{\mathcal{N}}{}$ :-$\coloneq\cfrac{\partial\mathbf{\tilde{u}}}{\partial t}+\mathbf{U}\cdot\nabla\mathbf{\tilde{u}}+\mathbf{\tilde{u}}\cdot\nabla\mathbf{U}+\mathbf{\tilde{u}}\cdot\nabla\mathbf{\tilde{u}}+\nabla\tilde{p}-\frac{1}{\text{Re}_{\tau}}\nabla^{2}{\mathbf{\tilde{u}}}=0,:$ (8)
@@
As mentioned above, $\langle{\mathbf{\tilde{u}}}\rangle_{xz}$ can be nonzero, and, as a consequence, $\langle{\mathbf{u}}\rangle_{xz}=\mathbf{U}+\langle{\mathbf{\tilde{u}}}\rangle_{xz}\neq\mathbf{U}$. Thus, it might appear like the initial disturbance $\mathbf{\tilde{u}}(\mathbf{x},0)$ is free to adjust $\langle{\mathbf{u}(\mathbf{x},0)}\rangle_{xz}$ to any arbitrary profile, in particular one that might be more suitable for generating disturbance energy growth than $\mathbf{U}$. However, the initial energies considered here are $\lesssim 0.02\%$ of the base energy (except for one run at $0.04\%$ in §4.3) and so this is not an issue.
The formulation given by Eqs. 8 and 9 is intended to be the simplest possible nonlinear model that could be used to investigate disturbance growth in turbulent flows. We note that in the literature, an alternative formulation exist (e.g. Farano et al., 2017), which differs from our work in that no background turbulence fluctuations are presumed. This, in turn, implies a feedback mechanism: Since in that model, the disturbances are responsible for maintaining the base flow $\mathbf{U}$, they are constrained to be in an order of magnitude that would be necessary for maintaining the mean flow in a DNS. For the questions to be investigated in the present work, this constrained behaviour of the disturbances would not make sense, as we are interested in how the disturbances grow on top of the base profile, drawing energy only from it.
Some authors considering the linear optimisation problem (Reynolds & Hussain, 1972; Pujals et al., 2009; Cossu & Hwang, 2017), have advocated for including an eddy viscosity in the momentum equation for the disturbances, i.e. 8. This introduces an ad-hoc assumption that the eddy viscosity felt by the mean profile is also the same as would be felt by the fluctuations. In order to keep the model as simple as possible, we decide not to include an eddy viscosity in this study, though its effect could be an interesting topic for future work.
@@
where $\lambda,\mathbf{\tilde{v}}$ and $\tilde{q}$ are Lagrange multipliers. Setting the first variation of $\mathcal{L}$ to zero leads, after some algebra, to the further conditions for the Lagrange multiplier fields that
@@
$\cfrac{\partial\mathbf{\tilde{v}}}{\partial t}+(\mathbf{U}+\mathbf{\tilde{u}})\cdot\nabla\mathbf{\tilde{v}}-\mathbf{\tilde{v}}\cdot{(\nabla(\mathbf{U}+\mathbf{\tilde{u}}))}^{\intercal}+\nabla\tilde{q}+\frac{1}{\text{Re}_{\tau}}\nabla^{2}{\mathbf{\tilde{v}}}$ $=0,$ (12)
@@
Note that Eqs. 13 and 12 form the adjoint Navier–Stokes equations, and the Lagrange multipliers $\mathbf{\tilde{v}}$ and $\tilde{q}$ can be interpreted as adjoint velocity and pressure (divided by the density), respectively. This set of equations naturally gives rise to a looping algorithm to obtain the optimal initial condition $\mathbf{\tilde{u}_{\text{opt}}}(\mathbf{x},0)$ that generates the highest energy gain $G$ at the end of the time horizon $T$: First, make an initial guess for $\mathbf{\tilde{u}_{\text{opt}}}(\mathbf{x},0)$, which we call $\mathbf{\tilde{u}}(\mathbf{x},0)$, and evolve it forward in time using the Navier–Stokes equations 9 and 8 until time $t=T$. Second, obtain the initial condition for the adjoint system $\mathbf{\tilde{v}}(\mathbf{x},T)$ using Eq. 11 and integrate Eqs. 13 and 12 backward in time from $t=T$ to $t=0$. Third, now that the final value of the adjoint variable $\mathbf{\tilde{v}}(\mathbf{x},0)$ is known, calculate ${\delta\mathcal{L}}/{\delta\mathbf{\tilde{u}}(\mathbf{x},0)}$ using Eq. 14 and use this gradient to improve the current guess $\mathbf{\tilde{u}}(\mathbf{x},0)$. Repeat this loop until some convergence criterion is reached. More technical details of this algorithm are discussed in Section 3.
@@ 3 Computational method
A special-purpose solver has been implemented in the programming language Python to integrate 9 and 8 forward in time and to integrate the adjoint system 13 and 12 backward in time. This solver utilises the library JAX (Bradbury et al., 2018), which makes the code fully differentiable (although this is not used much in the present work) and provides GPU-acceleration, thus delivering a crucial speed-up and making the investigation feasible. The source code of the solver is available at https://github.com/dakling/jax-spectral-dns.
@@
The effectiveness and efficiency of the looping approach crucially depends on the employed gradient-based optimisation algorithm. In the present work, we follow Cherubini et al. (2011) and use a conjugate gradient method, although we determine the step size adaptively using a two-way backtracking line search (Truong & Nguyen, 2020). This line search is similar to a regular backtracking line search, but the step size from the last iteration is used as an initial guess for the new step size, thus reducing the number of necessary function evaluations (which are quite expensive as they involve the forward run of the DNS). If the step size is found to be sufficiently small, it is also periodically tested if the step size can be increased, which is not necessary in classical backtracking line search. Note that the formula for the gradient 14 contains the Lagrange multiplier $\lambda$, which ensures that the initial energy constraint is fulfilled. Its appropriate value depends on the step size and direction taken, so the gradient itself depends on the parameters of the conjugate descent and must be recomputed appropriately. For a given set of descent parameters, $\lambda$ is then obtained using a standard Newton method. To ensure that numerical issues with determining $\lambda$ can never lead to a violation of the initial energy constraint, the resulting initial condition is also appropriately rescaled after each optimisation iteration. The algorithm is documented in detail in Section A.3.
@@ 4 Results
For any $T$, as $e_{0}\rightarrow 0$, the linear optimisation problem is recovered. Consequently, for small $e_{0}$, a quasilinear optimal, which closely resembles the linear optimal, is obtained. The linear optimal’s distinguishing property is that it consists of only a single wave number mode. This is because interactions between different wave number modes would require nonlinear effects, in the absence of which the optimisiation algorithm simply picks the wave number mode with the highest energy gain for the given time horizon. The quasilinear optimal is the weakly nonlinear extension of the linear optimal. Since it is finite amplitude, nonlinear coupling between modes exists but is too weak to make much of a difference compared to the linear optimal. As $e_{0}$ increases, there comes a point where a structurally completely different nonlinear optimal arises. For large $e_{0}$, the final state of the disturbances may be well in the turbulent regime, which, in the current formulation, usually leads the optimisation algorithm to fail, as the final state dependency on the initial condition is too sensitive and chaotic for the optimisation problem to remain well-conditioned. The precise determination of this boundary is difficult and not very important to the present study, but values up to $e_{0}/E_{0}=10^{-4}$ were found to converge without issues even for the longest time horizons considered here, where $E_{0}$ denotes the energy of the base flow. Naturally, we are most interested in nonlinear optimals whose initial energy is below this value.
@@
Figure 1: Optimal gain over initial energy normalised by the energy of the base flow $E_{0}$. Note the truncated vertical axis. Plots of the initial condition of the streamwise disturbance velocity for some of the optimals are also included. For more details on these and their time evolution please refer to Figs. 3 and 5.
[…]
Figure 2: Long-term evolution ($t_{\text{final}}u_{\tau}/h=60$) of the energy gain $G$ of the quasilinear optimal ($e_{0}/E_{0}=10^{-6},Tu_{\tau}/h=0.7$, blue line), the nonlinear optimal ($e_{0}/E_{0}={7.2}\cdot 10^{-5},Tu_{\tau}/h=0.7$, orange line), as well as the nonlinear optimal ($e_{0}/E_{0}={1}\cdot 10^{-4},Tu_{\tau}/h=2.8$, green line). Vertical dashed lines indicate $tu_{\tau}/h=0.7$ and $tu_{\tau}/h=2.8$, respectively. Snapshots of the streamwise disturbance velocity are shown for the ($e_{0}/E_{0}={7.2}\cdot 10^{-5}$, $Tu_{\tau}/h=0.7$)-optimal (orange line).
@@
Figure 4: Streamwise spectrum of the streamwise velocity at times $tu_{\tau}/h=0,0.35,0.7$ of the quasilinear optimal ($e_{0}/E_{0}=10^{-6}$) for $Tu_{\tau}/h=0.7$. Essentially all of the energy is contained in the $k_{x}=0$ mode.
Moving to finite but small values for $e_{0}/E_{0}$, at $10^{-6}$ and up to $10^{-5}$, a quasilinear optimal is found. Its velocity time evolution (not shown) closely resembles the linear optimal (Fig. 3), though the effects of the nonlinear term are visible in the combined energy of the non-dominant wave modes, which grows from around ${3}\cdot 10^{-4}$ to ${2.6}\cdot 10^{-3}$ times the total energy. This becomes apparent in the spectral decomposition of the streamwise velocity shown in Fig. 4.
Optimal gain is achieved by the nonlinear optimal with initial energy $e_{0}/E_{0}={7.2}\cdot 10^{-5}$, and its time evolution is shown in Figs. 5 and 6. It is apparent that the shape of the initial condition is completely different from the linear optimal. Most notably, no streaks are present, and the disturbance is concentrated, or localised, in specific parts of the domain. This localisation is most clear in spanwise direction, as there is no full localisation in the streamwise direction (see Fig. 6). Instead, in streamwise direction, there seems to be a length scale between the two distinct peaks of the disturbance. Indeed, when moving to a box twice as large in streamwise direction (i.e. $2\pi\times 2\times\pi$), four such peaks can be identified (see Section A.2). It is unclear how this length scale is determined, but we may speculate that it is the most efficient spacing to quickly generate continuous streaks. This may also be related to the size of the minimal channel unit, which is roughly half of the current domain in streamwise direction. Therefore, it is possible that the current structure could be thought of as two minimal channel optimals stitched together in streamwise direction. Localisation is also visible in the wall-normal direction, with the initial disturbance concentrated close to the wall where the base shear is high. The fact that the entire disturbance is concentrated on one side and not symmetrical can also be explained with localisation. This is in contrast to the linear optimal, for which both the symmetrical and the one-sided version yield the same gain, though we only show the one-sided version here. Localisation is a well-known phenomenon for nonlinear optimals (Kerswell, 2018), and can be rationalised by the optimisation algorithm trying to accommodate the initial energy constraint: instead of distributing the disturbance energy evenly, it may make more sense to have high-amplitude disturbances strategically placed in some small areas and then rely on nonlinear mechanisms to spread out these disturbances over time.
@@
Figure 9: Streamwise spectrum of the streamwise velocity at times $tu_{\tau}/h=0,0.35,0.7$ of the nonlinear optimal ($e_{0}/E_{0}={7}\cdot 10^{-5}$) for $Tu_{\tau}/h=0.7$. Compared to Fig. 4, the spectral dynamics are much more complicated, with streaks ($k_{x}=0$) not being dominant initially but then containing most of the energy later.
The early part of the evolution is characterised by an absence of streaks (see Fig. 9, left tile), and instead we observe tilted flow structures oriented such that they are amplified by an Orr-type mechanism (this is most clearly observed in Fig. 8). Following this initial phase, these disturbances quickly organise into streaks (see Figs. 6, 7 and 9, middle tile). Note that the streaks that have formed at this point do exhibit correct streak spacing and amplitude (see Fig. 5, middle tile, and the discussion in Section 4.2). Unlike the streaks of the linear optimal, the low-speed streaks do not remain straight as they evolve, but meander and wrap around the high-speed streaks, as is revealed by Fig. 7. Subsequently, these streaks grow due to the lift-up mechanism, and although no clear streak breakdown can be observed, it is clear from Fig. 10 that second-order disturbances grow on top of the streaks towards the end of the considered time window (from time (c) onwards). Thus, it can be concluded that the primary and, although not as pronounced, the secondary linear process are both visible in the nonlinear optimum, though the way that they are coupled together by the nonlinear term means that there is significant temporal overlap. This serves as a reminder that characterising them as neatly separated subprocesses constitutes a major simplification. Moreover, the streak growth itself is clearly not exclusively fuelled by transient growth (although this likely plays a significant role), as can be most clearly observed by the temporal evolution of the high-frequency modes ($k_{x}\geq 6$) between the times (a) and (b) shown in Fig. 10. After an initial growth up to the time marked by vertical line (a), these modes start decaying, indicating that they transfer energy to the streaks.